# Local-global compatibility and the exceptional zero conjecture for $\operatorname{GL}(3)$

### v1

@@ 9.7. The main results
Let $\mathfrak{m}_{0}\subset\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)$ denote a maximal ideal with residue field $k=\mathcal{O}/\varpi$, and suppose that the associated Galois representation $\overline{\rho}_{\mathfrak{m}_{0}}$ is absolutely irreducible and decomposed generic.
@@
    where $\vartheta\colon\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n},\mathcal{O}}\to\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)_{\mathfrak{m}_{0}}/J$ is the natural map.
@@
We first note that the Galois representation $\rho_{b,c,m}$ satisfying (1) already exists from the work of Scholze [Sch15] and Newton–Thorne [NT16] (see also [CGH+20]). Fix $1\leqslant j\leqslant d$, and let $\lambda$ be as in Lemma 9.6.10. We have a short exact sequence
$0\to\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},V_{\lambda,\mathcal{O}})^{\operatorname{ord}}/\varpi^{m}\to\operatorname{H}^{d-j}(X_{\operatorname{GL}_{n},K},\mathcal{O}/\varpi^{m})^{\operatorname{ord}}\to\operatorname{H}^{d-j+1}(X_{\operatorname{GL}_{n},K},V_{\lambda,\mathcal{O}})^{\operatorname{ord}}[\varpi^{m}]\to 0$
because $\mathcal{O}(\lambda)/\varpi^{m}\cong\mathcal{O}/\varpi^{m}$ and the ordinary part mod $\varpi^{m}$ only depends on the weight modulo $\varpi^{m}$. The result now follows from Proposition 9.6.12 (note that the kernel of the map
$\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(R\Gamma(X_{\operatorname{GL}_{n},K},\mathcal{O}/\varpi^{m})^{\operatorname{ord}})\to\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(\operatorname{H}^{*}(X_{\operatorname{GL}_{n},K},\mathcal{O}/\varpi^{m})^{\operatorname{ord}})$
is nilpotent, with nilpotence degree only depending on $n$). ∎
@@
For the last part, we follow [ACC+23, Lemma 6.2.11]. Namely, let $V$ denote the finite free of $L[\epsilon]$-module (of rank $n$) on which $\rho_{L[\epsilon]}$ acts. Define, iteratively, $W_{i}\subset V/W_{i-1}$ to be the largest submodule for which $\rho_{L[\epsilon]}$ acts through the character $f\circ\chi_{i}$ (with $W_{0}=0$). Let $V_{i}\subset V$ denote the preimage of $W_{i}$ under the map $V\to V/V_{i-1}$. Note that by property (3), we have $V_{n}=V$. We need to show that $V_{i}$ is free of rank $i$.
@@
$\operatorname{dim}_{L}(V/V_{n-1})/\epsilon=1$
(it cannot be zero, because this would imply $\prod_{i=1}^{n-1}(x-f\circ\chi_{i}(x))$ would kill $V$ modulo $\epsilon$ for all $x\in G_{\mathbb{Q}_{p}}$). By Nakayama’s lemma, we see that $V/V_{n-1}\neq 0$ and is free of rank one over $L[\epsilon]$ if we show it is $\epsilon$-torsion free.
@@
At the expense of shrinking $\Omega$, we may assume that $\Omega\subset\Omega^{\circ}\subset\Omega_{0}$, where $\Omega_{0}$ is a (sufficiently small) open affinoid neighbourhood of the trivial weight in $\mathcal{W}$ and $\Omega^{\circ}$ is a wide open disc, i.e., $\Omega^{\circ}$ is the adic generic fibre of $\operatorname{Spf}\mathcal{O}[\![T_{1}/p^{\rho},T_{2}/p^{\rho}]\!]$ for some $\rho\in\mathbb{Q}_{>0}$. Let $D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ denote the $\mathcal{O}^{+}(\Omega^{\circ})$-linear dual of $A_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ (the space of $r$-analytic (in the variables corresponding to $\operatorname{Iw}$) functions $f:\operatorname{Iw}\to\mathcal{O}^{+}(\Omega^{\circ})$ which satisfy $f(b\cdot-)=\kappa_{\Omega^{\circ}}(b)\cdot f(-)$ for all $b\in\overline{B}(\mathbb{Q}_{p})\cap\operatorname{Iw}$). The module $D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ carries actions of $(\operatorname{Iw},T^{+})$ via the same formulae as in §8.1.
@@
  • $\operatorname{Fil}^{0}D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}=D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ and $\bigcap_{i}\operatorname{Fil}^{i}D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}=0$;
@@
Let $F_{\Omega^{\circ}}\colon B(\mathbb{Z}_{p})\to T(\mathbb{Z}_{p})\xrightarrow{\kappa_{\Omega^{\circ}}}\Omega^{+}(\Omega^{\circ})$ denote the natural character, which we view as an element of $A_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ in the natural way. We have an $\mathcal{O}^{+}(\Omega^{\circ})$-linear map
@@
given by evaluating at $F_{\Omega^{\circ}}$. This map is equivariant for the action of $B(\mathbb{Z}_{p})$ as well as the action of $T^{+}$ (where $T^{+}$ acts via the projection $T^{+}\twoheadrightarrow T(\mathbb{Z}_{p})$ on the right-hand side). Let $\Lambda_{m}\colon\colon D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m}\to(\mathcal{O}^{+}(\Omega^{\circ})/\mathfrak{a}^{m})(\kappa_{\Omega^{\circ}}^{-1})$ denote the induced map, where $\mathfrak{a}$ denotes the maximal ideal of $\mathcal{O}^{+}(\Omega^{\circ})$. We have the following lemma:
@@
$R\Gamma(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}},$ $D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m})^{\operatorname{ord}}\cong R\Gamma(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}(0,m+r)},D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m})^{\operatorname{ord}}$
$\xrightarrow{\sim}R\Gamma(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}(0,m+r)},(\mathcal{O}^{+}(\Omega^{\circ})/\mathfrak{a}^{m})(\kappa_{\Omega^{\circ}}^{-1}))^{\operatorname{ord}}$
@@
which is the natural one away from $S$, and is given by $t\mapsto U_{t}\cdot\kappa_{\Omega^{\circ}}([t])$ for $t\in T^{+}$, factors through
@@
The morphism $\Lambda_{m}$ is equivariant for the action of $(\operatorname{Iw}(0,m+r),T^{+})$, so the map on cohomology complexes is well-defined. The first and third isomorphisms are standard using the fact that the ordinary part at level $\operatorname{Iw}(b,c)$ is independent of $c$ and [ACC+23, Lemma 5.2.9]. Finally, the second map is a quasi-isomorphism, because one can show that the action of $t=\operatorname{diag}(p^{2},p,1)\in T^{+}$ satisfies: $t\cdot\operatorname{ker}\Lambda_{m}\subset p\operatorname{ker}\Lambda_{m}$ (hence the ordinary part with coefficients in $\operatorname{ker}\Lambda_{m}$ vanishes). ∎
@@
given by identifying $\Sigma$ as an irreducible component of $\mathscr{E}\times_{\mathcal{W}}\Omega$ passing through the point $x_{\pi}$. Here $\mathscr{E}\to\mathcal{W}$ is the $\operatorname{GL}(3)$-eigenvariety in [Han17] constructed from (compactly-supported) overconvergent cohomology, and $x_{\pi}\in\mathscr{E}$ is the point corresponding to the unique $p$-stabillisation in $\pi$. By design, the map (9.8.2) factors $\underline{\alpha}^{S}\otimes\underline{\alpha}_{p}$, and the induced map $\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{3}}(\operatorname{H}_{c}^{*}(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}},D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}})^{\operatorname{ord}})\to\mathcal{O}(\Sigma)$ factors through the localisation at $\mathfrak{n}_{0}$ (because the specialisation of $\underline{\alpha}^{S}\otimes\underline{\alpha}_{p}$ at the trivial weight is the Hecke eigensystem associated with $\pi$).
@@
Note that, following this method, one can also obtain results about triangulations in families for ordinary points on the $\operatorname{GL}_{n}$-eigenvariety, for general $n\geqslant 2$. We leave the details to the interested reader.

### v2

@@ 9.7. The main results
Let $\mathfrak{m}_{0}\subset\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)$ denote a maximal ideal with residue field $k=\mathcal{O}/\varpi$, and suppose that the associated Galois representation $\overline{\rho}_{\mathfrak{m}_{0}}$ (see Theorem 9.2.7) is absolutely irreducible and decomposed generic.
@@
    where $\vartheta\colon\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{n}}\to\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)_{\mathfrak{m}_{0}}/J$ is the natural map.
@@
To simplify notation, in this proof only we write $X\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=X_{\operatorname{GL}_{n},K}$ and $\mathbf{T}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}$. By [KT17, Lem. 2.5(2)], the kernel $J_{1}$ of the map
$\theta_{1}\colon\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)=\mathbf{T}\Big(R\Gamma(X,\mathcal{O}/\varpi^{m})^{\operatorname{ord}}\Big)$ $\longrightarrow\oplus_{q=0}^{d-1}\mathbf{T}\Big(\operatorname{H}^{q}(X,\mathcal{O}/\varpi^{m})^{\operatorname{ord}}\Big)$
is nilpotent with nilpotence degree at most $d$ (hence only depending on $n$).
We first construct a representation over the target, working separately with each summand. Fix then $1\leqslant j\leqslant d$ (so q=djq=d-j), and let $\lambda$ be as in Lemma 9.6.10. We have a short exact sequence
$0\to\operatorname{H}^{d-j}(X,V_{\lambda,\mathcal{O}})^{\operatorname{ord}}/\varpi^{m}\to\operatorname{H}^{d-j}(X,\mathcal{O}/\varpi^{m})^{\operatorname{ord}}\to\operatorname{H}^{d-j+1}(X,V_{\lambda,\mathcal{O}})^{\operatorname{ord}}[\varpi^{m}]\to 0$
as $\mathcal{O}(\lambda)/\varpi^{m}\cong\mathcal{O}/\varpi^{m}$ and the ordinary part mod $\varpi^{m}$ only depends on the weight mod $\varpi^{m}$. Let
$\mathbf{T}(d-j,\lambda;m)$ . . $\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathbf{T}(\operatorname{H}^{d-j}(X,V_{\lambda,\mathcal{O}})^{\operatorname{ord}}/\varpi^{m}),$
$\mathbf{T}(d-j;m)$ . . $\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathbf{T}(\operatorname{H}^{d-j}(X,\mathcal{O}/\varpi^{m})^{\operatorname{ord}}),$
$\mathbf{T}^{\prime}(d-j+1,\lambda;m)$ . . $\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathbf{T}(\operatorname{H}^{d-j+1}(X,V_{\lambda,\mathcal{O}})^{\operatorname{ord}}[\varpi^{m}])$
be the respective Hecke algebras. From the short exact sequence, we obtain a surjective map
$\theta_{2}^{j}\colon\mathbf{T}(d-j;m)\longrightarrow\mathbf{T}(d-j,\lambda;m)\times\mathbf{T}^{\prime}(d-j+1,\lambda;m).$
Let $J_{2,j}=\ker(\theta_{2}^{j})$; a simple computation shows $T^{2}=0$ for all $T\in J_{2,j}$ (so $J_{2,j}$ is nilpotent).
Combining Proposition 9.6.12 with the natural maps of Hecke algebras, we have representations
$\varrho_{j}^{\prime}:G_{\mathbb{Q}}\to\operatorname{GL}_{n}(\mathbf{T}(d-j,\lambda;m)_{\mathfrak{m}_{0}}/J_{j}^{\prime}),\qquad\varrho^{\prime\prime}_{j}:G_{\mathbb{Q}}\to\operatorname{GL}_{n}(\mathbf{T}^{\prime}(d-j+1,\lambda;m)_{\mathfrak{m}_{0}}/J_{j}^{\prime\prime})$
satisfying the required conditions, where $J_{j}^{\prime}$ and $J_{j}^{\prime\prime}$ are nilpotent. Let $J_{3,j}=(\theta_{2}^{j})^{-1}(J_{j}^{\prime}\times J_{j}^{\prime\prime})$, which is nilpotent with nilpotence degree depending only on $n$ (noting $J_{2,j}\subset J_{3,j}$). Combining $\rho_{j}^{\prime}$ and $\rho_{j}^{\prime\prime}$ thus gives a representation
$\varrho_{j}:G_{\mathbb{Q}}\to\operatorname{GL}_{n}(\mathbf{T}(d-j;m)_{\mathfrak{m}_{0}}/J_{3,j})$
satisfying the required compatibilities. Let $\varrho=\oplus_{j=1}^{d}\varrho_{j}$.
Let $J\subset\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)_{\mathfrak{m}_{0}}$ be the kernel of the composition of $\theta_{1}$ (localised at $\mathfrak{m}_{0}$) with projection modulo $J_{3,1}\times\cdots\times J_{3,d}$. Note $J$, which contains $J_{1}$, is nilpotent of degree depending only on $n$, and $\theta_{1}$ induces an injective map
$\theta_{1}^{\prime}\colon\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)_{\mathfrak{m}_{0}}/J\hookrightarrow\oplus_{j=1}^{d}\mathbf{T}\big(d-j;m\big)_{\mathfrak{m}_{0}}/J_{3,j}.$
Now let $D_{\varrho}$ be the continuous determinant attached to $\varrho$, which satisfies the desired local-global compatibility (inherited from Proposition 9.6.12). By [Che14, Example 2.32], $D$ takes values in $\operatorname{Im}(\theta_{1}^{\prime})$, and pulling back under $\theta_{1}^{\prime}$ we obtain a continuous (Cayley–Hamilton) determinant $D\colon G_{\mathbb{Q}}\to\mathbf{T}_{\operatorname{GL}_{n}}^{S,\operatorname{ord}}(b,c;m)_{\mathfrak{m}_{0}}/J$ satisfying properties (1)–(3). We now obtain $\rho_{b,c,m}$ as the lift of $D$ under [Che14, Theorem 2.22(i)], using the fact that $\overline{\rho}_{\mathfrak{m}_{0}}$ is assumed to be absolutely irreducible. ∎
Remark 9.7.2.
The Galois representation $\rho_{b,c,m}$ satisfying (1) was previously constructed in the work of Scholze [Sch15] and Newton–Thorne [NT16] (see also [CGH+20]).
@@
For the last part, we follow [ACC+23, Lemma 6.2.11]. Namely, let $V$ denote the finite free $L[\epsilon]$-module (of rank $n$) on which $\rho_{L[\epsilon]}$ acts. Define, iteratively, $W_{i}\subset V/W_{i-1}$ to be the largest submodule for which $\rho_{L[\epsilon]}$ acts through the character $f\circ\chi_{i}$ (with $W_{0}=0$). Let $V_{i}\subset V$ denote the preimage of $W_{i}$ under the map $V\to V/V_{i-1}$. Note that by property (3), we have $V_{n}=V$. We need to show that $V_{i}$ is free of rank $i$.
@@
$\operatorname{dim}_{L}(V/V_{n-1})/\epsilon=1.$
By Nakayama’s lemma, we see that $V/V_{n-1}$ is cyclic over $L[\epsilon]$. We claim it is free of rank one; to see this, it suffices to show it is $\epsilon$-torsion free.
@@
At the expense of shrinking $\Omega$, we may assume that $\Omega\subset\Omega^{\circ}\subset\Omega_{0}$, where $\Omega_{0}$ is a (sufficiently small) open affinoid neighbourhood of the trivial weight in $\mathcal{W}$ and $\Omega^{\circ}$ is a wide open disc, i.e., $\Omega^{\circ}$ is the adic generic fibre of $\operatorname{Spf}\mathcal{O}[\![T_{1}/p^{\rho},T_{2}/p^{\rho}]\!]$ for some $\rho\in\mathbb{Q}_{>0}$.
Let $A_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ be the space of functions $f:\operatorname{Iw}\to\mathcal{O}^{+}(\Omega^{\circ})$ that are $r$-analytic in the variables corresponding to $\operatorname{Iw}$, and satisfy $f(b\cdot-)=\kappa_{\Omega^{\circ}}(b)\cdot f(-)$ for all $b\in\overline{B}(\mathbb{Q}_{p})\cap\operatorname{Iw}$. By restriction, we can identify $A_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ with the space of $r$-analytic functions $N(\mathbb{Z}_{p})\cong\mathbb{Z}_{p}^{3}\to\mathcal{O}^{+}(\Omega^{\circ})$. Let $D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ denote the $\mathcal{O}^{+}(\Omega^{\circ})$-linear dual of $A_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$. The module $D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ carries actions of $(\operatorname{Iw},T^{+})$ via the same formulae as in §8.1.
@@
  • $\operatorname{Fil}^{0}D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}=D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ and $\bigcap_{i}\operatorname{Fil}^{i}D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}=\{0\}$;
@@
Let $F_{\Omega^{\circ}}\colon B(\mathbb{Z}_{p})\to T(\mathbb{Z}_{p})\xrightarrow{\kappa_{\Omega^{\circ}}}\mathcal{O}^{+}(\Omega^{\circ})$ denote the natural character, which (for sufficiently large $r$) we view as an element of $A_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}$ in the natural way. Note that $F_{\Omega^{\circ}}$ corresponds to the constant function $1$ on $N(\mathbb{Z}_{p})$. We have an $\mathcal{O}^{+}(\Omega^{\circ})$-linear map
@@
given by evaluating at $F_{\Omega^{\circ}}$. This map is equivariant for the action of $B(\mathbb{Z}_{p})$ as well as the action of $T^{+}$ (where $T^{+}$ acts via the projection $T^{+}\twoheadrightarrow T(\mathbb{Z}_{p})$ on the right-hand side). Let $\Lambda_{m}\colon D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m}\to(\mathcal{O}^{+}(\Omega^{\circ})/\mathfrak{a}^{m})(\kappa_{\Omega^{\circ}}^{-1})$ denote the induced map, where $\mathfrak{a}$ denotes the maximal ideal of $\mathcal{O}^{+}(\Omega^{\circ})$. We have the following lemma:
@@
    $R\Gamma(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}},$ $D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m})^{\operatorname{ord}}\cong R\Gamma(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}(0,c)},D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m})^{\operatorname{ord}}$
    $\xrightarrow{\sim}R\Gamma(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}(0,c)},(\mathcal{O}^{+}(\Omega^{\circ})/\mathfrak{a}^{m})(\kappa_{\Omega^{\circ}}^{-1}))^{\operatorname{ord}}$ (9.8.2)
@@
    which is the natural one on $\mathbf{T}_{\operatorname{GL}_{3}}^{S}$ and given by $t\mapsto U_{t}\cdot\kappa_{\Omega^{\circ}}([t])$ for $t\in T^{+}$, factors through
@@
Part (ii) is a direct consequence of the Hecke-equivariance in (i). To see (i), first we explain why the claimed map exists. The first quasi-isomorphism is standard, using the fact that the ordinary part at level $\operatorname{Iw}(b,c)$ is independent of $c$ . The morphism $\Lambda_{m}$ is equivariant for the action of $(\operatorname{Iw}(0,c),T^{+})$, so the next map on cohomology complexes is well-defined. The third isomorphism is proved by first applying [ACC+23, Lemma 5.2.9] to pass from the cohomology of $X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}(0,c)}$ to that of $T(\mathbb{Z}_{p})$; then exploiting $R\Gamma(T(\mathbb{Z}_{p}),-)=R\Gamma(T(\mathbb{Z}_{p})/T(b),R\Gamma(T(b),-))$; then applying [ACC+23, Lemma 5.2.9] again to replace $R\Gamma(T(b),-)$ with the cohomology of $X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}(b,c)}$, pulling out $(\mathcal{O}^{+}(\Omega^{\circ})/\mathfrak{a}^{m})(\kappa_{\Omega^{\circ}}^{-1})$ (upon which $\operatorname{Iw}(b,c)$ acts trivially).
It remains to prove that the second map (9.8.2) is a quasi-isomorphism. We use the following:
Claim 9.8.3.
Let $t=\operatorname{diag}(p^{2},p,1)\in T^{+}$. We have $t^{r+1}\cdot\operatorname{ker}(\Lambda_{m})\subset p\operatorname{ker}(\Lambda_{m})$.
Given the claim, the $U_{p}$-operator acts on the cohomology with coefficients in $\operatorname{ker}\Lambda_{m}$ with strictly positive slope, and hence its ordinary part vanishes. The map (9.8.2) is thus a quasi-isomorphism from the long exact sequence attached to $0\to\ker(\Lambda_{m})\to D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}}/\operatorname{Fil}^{m}\to\mathcal{O}^{+}(\Omega^{\circ})/\mathfrak{a}^{m}\to 0$.
[…]
$(t^{-r-1}\cdot f)(X_{1},X_{2},X_{3})=a_{0}+g(pX_{1},p^{2}X_{2},pX_{3}).$
In particular, as $g$ has no constant term, we can write $t^{-r-1}\cdot f=a_{0}+pg^{\prime}$, for $g^{\prime}\in\mathcal{O}^{+}(\Omega^{\circ})\langle X_{1},X_{2},X_{3}\rangle$ with $g^{\prime}(0,0,0)=0$. Note the association $f\mapsto g^{\prime}$ is $\mathcal{O}^{+}(\Omega^{\circ})$-linear.
Now let $\mu]\in\ker(\Lambda_{m})$, represented by some $\mu$ in the dual of $r$-analytic functions on $N(\mathbb{Z}_{p}$); then $\mu(a_{0})=a_{0}\mu(F_{\Omega^{\circ}})\equiv 0\hskip 2.0pt(\mathrm{mod}\hskip 2.0pt\mathfrak{a}^{m})$ for any constant function $a_{0}\in\mathcal{O}^{+}(\Omega^{\circ})$ (recalling $F_{\Omega^{\circ}}$ corresponds to the constant function $1$). Define $\lambda\in D^{\circ,r\operatorname{-an}}_{\Omega^{\circ}}$ by $\lambda(f)=\mu(g^{\prime})$, for notation as above. If $f=F_{\Omega^{\circ}}$, then $g^{\prime}=0$; so $\lambda(F_{\Omega^{\circ}})=\mu(0)=0$, and hence $\lambda]\in\ker(\Lambda_{m})$. We find
$(t^{r+1}\cdot\mu)(f)=\mu(t^{-r-1}\cdot f)=\mu(a_{0}+pg^{\prime})\equiv p\lambda(f)\hskip 2.0pt(\mathrm{mod}\hskip 2.0pt\mathfrak{a}^{m}).$
We see $t^{-r-1}\cdot[\mu]=[t^{-r-1}\cdot\mu]=p[\lambda]\in p\ker(\Lambda_{m})$, as required. ∎
@@
given by identifying $\Sigma$ as an irreducible component of $\mathscr{E}\times_{\mathcal{W}}\Omega$ passing through the point $x_{\pi}$. Here $\mathscr{E}\to\mathcal{W}$ is the $\operatorname{GL}(3)$-eigenvariety in [Han17] constructed from (compactly-supported) overconvergent cohomology, and $x_{\pi}\in\mathscr{E}$ is the point corresponding to the unique $p$-stabilisation in $\pi$. By design, the map (9.8.4) factors $\underline{\alpha}^{S}\otimes\underline{\alpha}_{p}$, and the induced map $\mathbf{T}^{S,\operatorname{ord}}_{\operatorname{GL}_{3}}(\operatorname{H}_{c}^{*}(X_{\operatorname{GL}_{3},K^{p}\operatorname{Iw}},D_{\Omega^{\circ}}^{\circ,r\operatorname{-an}})^{\operatorname{ord}})\to\mathcal{O}(\Sigma)$ factors through the localisation at $\mathfrak{n}_{0}$ (because the specialisation of $\underline{\alpha}^{S}\otimes\underline{\alpha}_{p}$ at the trivial weight is the Hecke eigensystem associated with $\pi$).
@@
We have specialised to $\operatorname{GL}_{3}$ for our intended application; but following this method, one can also obtain results about triangulations in families for ordinary points on the $\operatorname{GL}_{n}$-eigenvariety, for general $n\geqslant 2$.